The sample space of an experiment is:
{A, B, C, D} AND P(A) = 0.1, P(B) = 0.3, P(C) = 0.4, P({A,C})=
: The sample space of an experiment is:
{A, B, C, D} AND P(A) = 0.1, P(B) = 0.3, P(C) = 0.4, P({A,C})=

Answer by stanbon(19693)

The sample space of an experiment is:
{A, B, C, D} AND P(A) = 0.1, P(B) = 0.3, P(C) = 0.4, P({A,C})=
---------------
IF A and C are independent events the following is true:
P(A and C) = P(A)*P(C) = 0.1*0.4 = 0.04
----------------
If A and C are dependent events the following is true:
P(A and C) = P(A)*P(C|A) = 0.1*? = ?
----------------
Cheers,
Stan H.